OpenStudy (elifbestegul):

Determine the refractive index of a certain type of glass in which the velocity of light is 2.40 x 10^8 m.s^-1

OpenStudy (nikks1992):

as you know RI of any medium = c/v ,where c= velocity of light in air and v= velocity of light in medium applying it here RI = 3*10^8/2.40 x 10^8 = 1.25

OpenStudy (elifbestegul):

thank you so much again @nikks1992

OpenStudy (nikks1992):

Well keep asking :)